### v1

# Dimensional Uplift in Conformal Field Theories

Abstract
The n-point functions of any Conformal Field Theory (CFT) in $d$ dimensions can always be interpreted as spatial restrictions of corresponding functions in a higher-dimensional CFT with dimension $d^{\prime}>d$. In particular, when a four-point function in $d$ dimensions has a known conformal block expansion, this expansion can be easily extended to $d^{\prime}=d+2$ due to a remarkable identity among conformal blocks, discovered by Kaviraj, Rychkov, and Trevisani (KRT) as a consequence of Parisi-Sourlas supersymmetry and confirmed to hold in any CFT with $d>1$.
In this note, we provide an elementary proof of this identity using simple algebraic properties of the Casimir operators. Additionally, we construct five differential operators, $\Lambda_{i}$, which promote a conformal block in $d$ dimensions to five conformal blocks in $d+2$ dimensions. These operators can be normalized such that $\sum_{i}\Lambda_{i}=1$, from which the KRT identity immediately follows. Similar, simpler identities have been proposed, all of which can be reformulated in the same way.
Dimensional Uplift in Conformal Field Theories
Ferdinando Gliozzi
Dipartimento di Fisica dell’Università di Torino
Regge Center for Algebra, Geometry and Theoretical Physics,
via Pietro Giuria  1, I-10125 Torino, Italy
1 Introduction
Dimensional reduction is a property emerging in some disordered systems. In the random-field Ising model it was observed that the correlation functions reduced to those of the pure Ising model, without the random source, in two fewer dimensions [1, 2]. This was understood by Parisi and Sourlas [3] to be due to a hidden supersymmetric formulation. This dimensional reduction holds near the upper critical dimension of six, but fails in sufficiently low dimensions [4, 5, 6, 7, 8, 9, 10, 11, 12]. This failure is now believed to result from operators that are irrelevant near six dimensions becoming relevant around five dimensions [6, 10, 13].
Another notable example is the connection, again discovered by Parisi and Sourlas [14], between randomly branched polymers or lattice animals in $d$ dimensions [15, 16, 17, 18, 19] and the Yang-Lee edge singularity [20] in two fewer dimensions. In this case, dimensional reduction works not only near the upper critical dimension of eight but continues down to two dimensions, as rigorously demonstrated by Brydges and Imbrie [21, 22]. Even in these instances, the correlation functions on branched polymers reduce to those of the Yang-Lee edge singularity.
In the present note we explore such a connection across the space dimensions from the point of view of conformal field theory (CFT). As a result, we give a simple algebraic proof of a surprising exact relation discovered by Kaviraj, Rychkov and Trevisani (KRT) as a direct consequence of the Parisi-Sourlas (PS) supersymmetry and confirmed to hold in any CFT with $d>1$ [10]; this relation expresses any $(d-2)$-dimensional conformal block as a linear combination of five conformal blocks in dimension $d$. We in a sense reverse such a relation by constructing, only utilizing some commutation properties of the Casimir operators, five differential operators, $\Lambda_{i}$, transforming an arbitrary $(d-2)$-dimensional conformal block $g^{(d-2)}_{\Delta,\ell}$ ( describing the contribution to a four-point function of a primary of scaling dimension $\Delta$ and spin $\ell$) into five different conformal blocks in dimenion $d$,
$\Lambda_{i}\,\,g^{(d-2)}_{\Delta,\ell}\propto\,g^{(d)}_{\Delta_{i},\ell_{i}}\,,$ (1)
where the conformal blocks in the rhs are precisely those contributing to the KRT identity. Equation (1) can be utilized to build new conformal blocks in higher dimensions once their form in lower dimensions is known. The $\Lambda_{i}$’s can be normalized so that
$\sum_{i=1}^{5}\Lambda_{i}=1~{}~{},$ (2)
therefore $\sum_{i}\Lambda_{i}\,g^{(d-2)}_{\Delta,\ell}$ yields at once the KRT identity. In such a form it is clear that this identity holds for any real $d>1$ under the only assumption that the conformal blocks are eigenfunctions of the Casimir operators.
The KRT identity allows one to reconstruct the spectrum of primary operators contributing to a four-point function of a CFT in dimension $d$ once this spectrum is known in a $(d-2)$-dimensional CFT with the same external scalars [23].  [^1] In [23] this result is obtained by assuming the PS supersymmetry.
2 Notation
In a generic CFT in $d$ dimensions the 4pt function of arbitrary scalars $\mathcal{O}_{i}$ of scaling dimension $\Delta_{i}$ can be parametrized as [24, 25]
$\langle\mathcal{O}_{1}(x_{1})\mathcal{O}_{2}(x_{2})\mathcal{O}_{3}(x_{3})\mathcal{O}_{4}(x_{4})\rangle=K(x_{i})\,g(u,v)\,,$ (3)
where $K(x_{i})$ is a kinematic factor given by
$K(x_{i})=\frac{1}{|x_{12}|^{\Delta_{1}+\Delta_{2}}\,|x_{34}|^{\Delta_{3}+\Delta_{4}}}\,\left(\frac{x_{14}^{2}}{x_{24}^{2}}\right)^{a}\,\left(\frac{x_{14}^{2}}{x_{13}^{2}}\right)^{b}\,,$ (4)
with $a=\frac{\Delta_{2}-\Delta_{1}}{2}$ and $b=\frac{\Delta_{3}-\Delta_{4}}{2}$ ; $g(u,v)$ is a theory-dependent function of the two cross-ratios $u=\frac{x_{12}^{2}x_{34}^{2}}{x_{13}^{2}x_{24}^{2}}$ and $v=\frac{x_{14}^{2}x_{23}^{2}}{x_{13}^{2}x_{24}^{2}}$, which are linearly independent for $d>1$.
It is important to emphasize that although $g(u,v)$ depends on the scaling dimensions of the external scalars, it is completely independent of the space-time dimension $d$:
$\partial_{d}\,\,g(u,v)=0\,\,$ (5)
In principle, we can choose $d$ to be any real number, provided that $d\geq\sf d$, where $\sf d$ represents the dimension of the linear space spanned by the four points $x_{i}$ (hence, $1<\sf d\leq 3$). Under this condition, we obtain the expansion
$g(u,v)=\sum_{\Delta,\ell}\,c_{\Delta,\,\ell}\,\,g^{(d)}_{\Delta,\ell}(u,v),$ (6)
where the coefficients $c_{\Delta,\,\ell}$ are the operator product expansion (OPE) coefficients.
The freedom in choosing $d$ has intriguing physical consequences. For instance, if $g(u,v)$ corresponds to the four-point function of a random branched polymer system in dimension $d$, then, due to dimensional reduction, its expansion in $(d-2)$-dimensional conformal blocks provides insight into the spectrum of primary operators associated with the Yang-Lee edge singularity. The dimensional shift by 2 in this case is well understood in terms of Parisi-Sourlas supersymmetry.
A natural question then arises: can this reduction by 2, or the reverse approach, the dimensional uplift, be explained purely through conformal invariance? A simple example helps to clarify the issue. Consider the four-point function $\langle\phi(x_{1})\phi(x_{2})\phi(x_{3})\phi(x_{4})\rangle\,$, where $\phi$ is a scalar generalized free field (GFF) with an arbitrary scaling dimension $\delta$. The corresponding function $g(u,v)$ is given by
$g(u,v)=1+u^{\delta}+\left(\frac{u}{v}\right)^{\delta}.$ (7)
The OPE coefficients and the spectrum of contributing primaries have been exactly evaluated in various ways [26, 27, 28]. Since $\delta$ is an arbitrary real number, generalized free CFTs are not necessarily local CFTs. For locality, the theory must have a non-vanishing coupling to the conserved energy-momentum tensor $T_{\mu\,\nu}$, which is a primary operator with scaling dimension $d$ and spin 2. The spectrum of primaries contributing to the expansion of $g(u,v)$ is
$\{\Delta,\ell\}=\{2\delta+2m+2n,\ell=2n\},\quad(m,n=0,1,2,\dots)\,.$ (8)
Keeping $\delta$ fixed, if the equation $2\delta+2m+2=d$ has a solution, then clearly there exist infinitely many solutions for dimensions $d^{\prime}=d+2k$ with $k=1,2,\dots$
This behaviour seems to be a general feature of local conformal field theories. Actually, as already pointed out in [10], the KRT identity implies that if a scalar primary of scaling dimension $\delta$ has a non-vanishing coupling to the energy-momentum tensor in dimension $d$, then it describes a local CFT in any space dimension $d^{\prime}=d+2k\,\,$ , $(k\in{\mathbb{N}})$ , and this property holds for any real $d>1$.
3 The proof
The quadratic Casimir operator $C_{2}[d]$ can be written as [24]
$C_{2}[d]=D_{z}+D_{\bar{z}}+(d-2)\frac{z\,{\bar{z}}}{z-{\bar{z}}}\left((1-z)\partial_{z}-(1-{\bar{z}})\partial_{\bar{z}}\right)\,,$ (9)
with
$D_{z}=z^{2}(1-z)\partial_{z}^{2}-(1+a+b)z^{2}\partial_{z}+ab\,z~{}.$ (10)
The two variables $z,{\bar{z}}$ are related to the cross-ratios as $u=z\,{\bar{z}}$ and $v=(1-z)(1-{\bar{z}})$.
We aim to demonstrate a special relationship between the quadratic Casimir oprators in $d$ and $d-2$ dimensions. Specifically, we iteratively construct five differential operators $Y_{i}$ $(i=1,\dots 5)$ where $Y_{1}=1$, $Y_{2}$ is first-order in $z$ and ${\bar{z}}$, $Y_{3}$ and $Y_{4}$ are second-order, and $Y_{5}$ is third-order, such that
$C_{2}[d]\,Y_{i}=Y_{i}\,C_{2}[d-2]+Y_{j}\,\alpha^{j}_{i}\,\,\,,$ (11)
where the $\alpha^{j}_{i}$’s are simple linear functions of the Casimir in dimension $d-2$. If we attempt to replace in the rhs $C_{2}[d-2]$ with $C_{2}[d^{\prime}]$ with $d^{\prime}\not=d-2$, the linearity in the $Y_{i}$ is immediately lost, and the algebraic structure of the equation no longer closes.
Basic linear algebra manipulations of this equation allow us to define five operators $\Lambda_{i}$, which promote any conformal block in $d-2$ dimensions to a suitable conformal block in $d$ dimensions - leading to the KRT identity.
We begin with $Y_{1}=1$ and write
$C_{2}[d]\,Y_{1}=Y_{1}\,C_{2}[d-2]+Y_{2}~{},$ (12)
where $Y_{2}=\frac{2z\,{\bar{z}}}{z-{\bar{z}}}\left((1-z)\,\partial_{z}-(1-{\bar{z}})\,\partial_{\bar{z}}\right)$.   We proceed with
$C_{2}[d]\,Y_{2}=Y_{2}\,(C_{2}[d-2]-d+2)+2\,Y_{1}\,C_{2}[d-2]+2\,Y_{3}-2\,Y_{4}~{},$ (13)
where $Y_{3}=\frac{z\,{\bar{z}}}{z-{\bar{z}}}(D_{z}-D_{\bar{z}})$ and $Y_{4}=\frac{z+{\bar{z}}}{z{\bar{z}}}\,Y_{3}$.
Now we have
$C_{2}[d]\,Y_{3}=Y_{3}\,(C_{2}[d-2]-d+2)~{};$ (14)
this remarkable relation, first discovered by Dolan and Osborn [24] and recently reobtained by Trevisani [23] as a consequence of the PS supersymmetry, tells us that $Y_{3}\,g^{(d-2)}_{\Delta,\ell}$ is proportional to $g^{(d)}_{\Delta+1,\ell-1}$.
Next we obtain
$C_{2}[d]\,Y_{4}=Y_{4}\,(C_{2}[d-2]+2)+Y_{3}\,(2(a+b)-d)+2\,Y_{5}~{},$ (15)
where $Y_{5}=((z-1)\,\partial_{z}+({\bar{z}}-1)\,\partial_{\bar{z}})Y_{3}$.
Finally we find
$C_{2}[d]\,Y_{5}=$ $Y_{5}\,(C_{2}[d-2]-3d+8)+Y_{4}\,(C_{2}[d-2]-(d-2)^{2})-Y_{1}\,C_{4}[d-2]$ (16)
$+Y_{3}\,(2ab-2(d-3)(a+b)+(d-2)^{2}-C_{2}[d-2])~{},$
where $C_{4}$ is the quartic Casimir operator defined by Dolan and Osborn in [24] :
$C_{4}[d]=\left(\frac{z{\bar{z}}}{z-{\bar{z}}}\right)^{d-2}\left(D_{z}-D_{\bar{z}}\right)\left(\frac{z{\bar{z}}}{z-{\bar{z}}}\right)^{-d+1}\,Y_{3}~{}.$ (17)
Clearly, when we apply the $C_{2}[d]\,Y_{i}$’s to an arbitrary conformal block in $d-2$ dimensions, these differential operators span a five-dimensional space where the action of $C_{2}[d]$ is represented by a $5\times 5$ matrix:
$C_{2}[d]{\sf Y}=\textstyle\left(\begin{matrix}c_{2}&1&0&0&0\cr 2c_{2}&c_{2}-d+2&2&-2&0\cr 0&0&c_{2}-d+2&0&0\cr 0&0&2(a+b)-d&c_{2}+2&2\cr-c_{4}&0&\kappa&c_{2}-(d-2)^{2}&c_{2}-3d+8\cr\end{matrix}\right){\sf Y}$ (18)
where ${\sf Y}$ is the column-vector formed by the $Y_{i}$’s, $\kappa$ is the coefficient of $Y_{3}$ in Eq.(16) and $c_{2},c_{4}$ are the eigenvalues of the quadratic and quartic Casimir operators in $d-2$ dimensions:
$c_{2}=$ $\ell(\ell+d-4)/2+\Delta(\Delta-d+2)/2\,\,\,;$ (19)
$c_{4}=$ $-\ell(\ell+d-4)(d-\Delta-3)(\Delta-1)\,\,\,\,.$ (20)
One can check at once that the five eigenvectors $V_{i},(i=1,2,..,5)$ of the transposed of such a matrix define five differential operators $\Lambda_{i}=V_{i}\cdot Y$ which transform $g^{(d-2)}_{\Delta,\ell}$ into a suitable conformal block in $d$ dimensions.
$\Lambda_{i}\,g^{(d-2)}_{\Delta,\ell}\equiv\Lambda_{\alpha\,\beta}\,\,g^{(d-2)}_{\Delta,\ell}\propto g^{(d)}_{\Delta+\alpha,\ell-\beta}\,,\,\alpha,\beta\in\{0,1,2\},\,\alpha+\beta\,\,\,{\rm even}$ (21)
Precisely we obtain
$\Lambda_{00}$ $=\textstyle~{}n_{00}\,\big{[}Y_{5}+(3d-8+\ell-\Delta)\,Y_{4}+(a+b+\frac{2a\,b}{d-2+\ell-\Delta}-\frac{3d-8+\ell-\Delta}{2})\,Y_{3}+$
$\textstyle~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\frac{(d-4+\ell)(3-d+\Delta)}{2}\,Y_{2}+\frac{(d-4+\ell)(3-d+\Delta)(d+\ell-\Delta+2)}{2}\big{]}\,,$
$\Lambda_{20}$ $=\textstyle~{}n_{20}\big{[}Y_{5}+\frac{2d-6+\ell-\Delta}{2}\,Y_{4}+(3+a+b+\frac{2a\,b}{\ell+\Delta}-\frac{\Delta+\ell}{2})\,Y_{3}-$
$\textstyle~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-\frac{(\Delta-1)(d-4+\ell)}{2}\,Y_{2}-\frac{(d-4+\ell)(\Delta-1)(\Delta+\ell)}{2}\big{]}\,,$
$\Lambda_{11}$ $=\textstyle~{}n_{11}\,Y_{3}\,,$ (22)
$\Lambda_{02}$ $=\textstyle~{}n_{02}\big{[}Y_{5}+(d-2-\frac{\ell+\Delta}{2})\,Y_{4}+(a+b-\frac{2a\,b}{\ell+\Delta-2}+\frac{4-2d+\ell+\Delta}{2}+\ell+\Delta-2)\,Y_{3}+$
$\textstyle~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\frac{\ell(d-3-\Delta)}{2}\,Y_{2}+\frac{\ell(d-3-\Delta)(\Delta+\ell-2)}{2}\big{]}\,,$
$\Lambda_{22}$ $=\textstyle~{}n_{22}\big{[}Y_{5}+\frac{d-2-\ell+\Delta}{2}\,Y_{4}+\frac{2(a+b+1)+\ell-\Delta-d-\frac{4a\,b}{d-4+\ell-\Delta}}{2}\,Y_{3}+$
$\textstyle~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\frac{\ell(\Delta-1)}{2}\,Y_{2}-\frac{\ell(\Delta-1)(d-4+\ell-\Delta)}{2}\big{]}\,.$
The normalization coefficients $n_{\alpha\,\beta}$ can be chosen in such a way that $\sum_{i}\Lambda_{i}=1$, which implies of course five linear equations for the $n_{\alpha\,\beta}$’s. Thier solution yields
$n_{00}$ $=\textstyle~{}\frac{-2}{(d-4+2\ell)(d-2\Delta)(d-3+\ell-\Delta)}\,,$
$n_{20}$ $=\textstyle~{}\frac{-2}{(d-4+2\ell)(d-2-2\Delta)(\ell+\Delta-1)}\,,$
$n_{11}$ $=\textstyle~{}\frac{-8a\,b}{(d-4+\ell-\Delta)(d-2+\ell-\Delta)(\ell+\Delta-2)(\ell+\Delta)}\,,$ (23)
$n_{02}$ $=\textstyle~{}\frac{2}{(d-4+2\ell)(d-2-2\Delta)(d-3+\ell-\Delta)}\,,$
$n_{22}$ $=\textstyle~{}\frac{2}{(d-4+2\ell)(d-2-2\Delta)(\ell+\Delta-1)}\,.$
Taking advantage of this normalization we finally obtain the KRT identyty
$g^{(d-2)}_{\Delta,\ell}\equiv\sum_{i=1}^{5}\Lambda_{i}\,g^{(d-2)}_{\Delta.\ell}=\sum_{\alpha,\beta\in\{0,1,2\},\,\alpha+\beta\,\,{\rm even}}k_{\alpha\,\beta}\,g^{(d)}_{\Delta+\alpha,\ell-\beta}~{}.$ (24)
The $k_{\alpha\,\beta}$’s depend of course on the normalization of the conformal blocks. In accordance with [10] we have
$k_{20}$ $=\textstyle-\frac{(\Delta-1)\Delta\left(\Delta-\Delta_{12}+\ell\right)\left(\Delta+\Delta_{12}+\ell\right)\left(\Delta-\Delta_{34}+\ell\right)\left(\Delta+\Delta_{34}+\ell\right)}{4(d-2\Delta-4)(d-2\Delta-2)(\Delta+\ell-1)(\Delta+\ell)^{2}(\Delta+\ell+1)}\,,$
$k_{11}$ $=\textstyle-\frac{(\Delta-1)\Delta_{12}\Delta_{34}\ell}{(\Delta+\ell-2)(\Delta+\ell)(d-\Delta+\ell-4)(d-\Delta+\ell-2)}\,,$ (25)
$k_{02}$ $=\textstyle-\frac{(\ell-1)\ell}{(d+2\ell-6)(d+2\ell-4)},\,~{}~{}~{}k_{00}=1,$
$k_{22}$ $=\textstyle\frac{(\Delta-1)\Delta(\ell-1)\ell\left(d-\Delta-\Delta_{12}+\ell-4\right)\left(d-\Delta+\Delta_{12}+\ell-4\right)\left(d-\Delta-\Delta_{34}+\ell-4\right)\left(d-\Delta+\Delta_{34}+\ell-4\right)}{4(d-2\Delta-4)(d-2\Delta-2)(d+2\ell-6)(d+2\ell-4)(d-\Delta+\ell-5)(d-\Delta+\ell-4)^{2}(d-\Delta+\ell-3)}\,.$
Similar, simpler dimensional uplift identities for two-point and three-point functions of CFT with boundary [29, 30] and for two-point functions in real projective space [31] have been found. All of them can be reformulated in terms of differential operators transforming a conformal block in dimension $d-2$ into a conformal block in dimension $d$. For instance, in the case of CFT with boundaries, the boundary conformal blocks $f_{bdy}^{(d)}[\Delta,\xi]$ are eigenfunctions of the following quadratic Casimir
${\cal C}[d]\,=\,-\xi(\xi+1)\partial_{\xi}^{2}-d(\xi+\frac{1}{2})\partial_{\xi}~{},~{}{\cal C}[d]\,f^{(d)}_{bdy}[\Delta,\xi]=\,\Delta\,(d-1-\Delta)\,f_{bdy}^{(d)}[\Delta,\xi]$ (26)
This operator acts linearly on $\{y_{1}=1,y_{2}=-(1+2\xi)\,\partial_{\xi}\}$, namely
${\cal C}[d]\,y_{1}=\,y_{1}\,{\cal C}[d-2]+y_{2}$ (27)
${\cal C}[d]\,y_{2}=\,y_{2}\,({\cal C}[d-2]+2(d-3))-4y_{1}\,{\cal C}[d-2]\,.$ (28)
By applying the same approach as before we build two differential operators, $\lambda_{0}$ and $\lambda_{2}$,
$\lambda_{0}=\frac{y_{2}+2(\Delta-d+3)}{2(2\Delta-d+3)}\,,~{}~{}\lambda_{2}=\frac{-y_{2}+2\Delta}{2(2\Delta-d+3)}\,\,\,,~{}~{}~{}\lambda_{0}+\lambda_{2}=1\,\,\,.$ (29)
such that  [^2] Since it has been noticed [32] that the three-dimensional boundary conformal blocks are expressible as elementary algebraic functions, the mere existence of $\lambda_{0}$ and $\lambda_{2}$ allows us to conclude that this property holds for any odd $d$.
$\lambda_{0}\,f_{bdy}^{((d-2)}[\Delta,\xi]=f_{bdy}^{(d)}[\Delta,\xi]\,~{}~{},\,\lambda_{2}\,f_{bdy}^{(d-2)}[\Delta,\xi]=k\,f_{bdy}^{(d)}[\Delta+2,\xi]\,\,,$ (30)
[…]
4 Conclusions
We provided a straightforward algebraic proof of a remarkable identity discovered some time ago by Kaviraj, Rychkov, and Trevisani (KRT). This identity expresses a conformal block contributing to the four-point function in $(d-2)$ dimensions as a linear combination of five conformal blocks in $d$ dimensions. The key mechanism involved the iterative construction of five differential operators that transform linearly under the action of the quadratic Casimir. This, in turn, enabled us to define five differential operators, denoted by $\Lambda_{i}$, which transform any conformal block in $(d-2)$ dimensions into five conformal blocks in $d$ dimensions. These operators can be normalized such that $\sum_{i}\,\Lambda_{i}=1$, from which the KRT identity follows immediately. Similar, simpler identities have been proposed in recent literature, all of which can be reformulated using the present approach. As an example, we explicitly worked out the case of boundary conformal blocks contributing to the two-point functions of boundary conformal field theories.